OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (cwrw238):

x + 4y = 6 y = (-1/4)x + 3/2 slope of the line parallel = -1/4 using y - y1 = m (x - x1) y - 5 = (-1/4) ( x - -8) y = (-1/4)(x + 8) + 5 y = (-1/4)x + 3